# Mixture of Cluster-conditional LoRA Experts for Vision-language Instruction Tuning

### 08 Feb 2024 v3

@@ 4.5 Visualizations
Figure 4: Clustering assignment of the training datasets when $K=64$. The labels on the vertical axis indicate the names of the datasets and they are arranged such that different datasets of the same task are placed together. Within each labeled block, multiple rows represent different subsets, each of which is associated with a unique task template. The horizontal axis denotes the cluster index to which these subsets are assigned. Darker color indicates more samples of this subset are assigned to certain cluster.
Clustering.
@@
We first show the justification to represent the training data via their instructions. Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model [43]. After that, we visualize the data in Figure 3 via t-SNE [51]. According to Figure 3, we observe that (1) samples from different datasets but from the same task are grouped together. For example, coco_vqg, aok_vqg, and ok_vqg are all VQG (visual question generation) tasks (triangle markers) and reside on the left part of the figure. (2) samples from similar tasks are closed to each other. For example, coco_cap and textcaps both belong to the image captioning (CAP) task (small dots) and stay close to each other. Similarly, both the visual question answering (VQA, cross markers) task with datasets (e.g., aok_vqa, coco_vqa, ocrvqa, and ok_vqa) and conversation (CONV) task with datasets (e.g., llava_conversation, llava_detail, and llava_reason) involve answering user questions, which lie in the middle part of Figure 3, suggesting that instructions are good representatives of the training data.
@@
We then cluster all the instructions of the examples in the training data into 64 groups using $k$-means clustering. Figure 4 shows the cluster assignment of the training data. Here, each row in the heatmap denotes a subset of a dataset. The subset is obtained by applying the task template (Sec. 3.1) on the samples of the dataset. We observe (1) different subsets of the same datasets are assigned to similar clusters. For example, aok_vqa, coco_vqa, and ok_vqa are in the first several clusters, while aok_vqg, coco_vqg, and ok_vqg mainly occupy 4 other clusters. (2) datasets of similar tasks are assigned to common clusters. For example, llava_150k including llava_detail, llava_detail and llava_conversation and a series of VQA tasks share the first several clusters as they are to answer questions. Those evidences justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
Routing results.
@@
Figure 5 visualizes the routing decisions of the proposed MoCLE and Sentence-MoLE. We obtain both results from one mixture of LoRA, i.e., one linear module in a layer. The routing results are aggregated by the subset of datasets similar to Figure 4. Shown in Figure 4(a), MoCLE could achieve task-level routing for the inputs. For example, datasets from VQA and VQG tasks are handled by expert 0 and 3 respectively. In addition, VQA expert (i.e., expert 0, which is mostly occupied by VQA task) is fundamental for most of the datasets. That is, most held-out datasets activate the VQA expert. This could be justified by the fact that most vision-language tasks take the form of question answering. However, this does not indicate that the entire model collapses to a single expert because knowledge from other tasks are learned by the universal expert and are leveraged during evaluation. Further, the remaining 3 experts are not dispensable as without them all tasks would clutter in a single expert, resulting in task conflicts.
On the contrary, routing pattern of Sentence-MoLE in Figure 4(b) reveals little correlations between tasks/datasets and experts. That is, different datasets/tasks obtain similar routing decisions, still suffering from task conflicts.
@@
Figure 5: Routing decisions of one LoRA mixture for MoCLE and Sentence-MoLE. The setup of the vertical axis is similar to Fig. 4 except that we also include the held-out tasks. They are separated by a dotted line on the vertical axis. The horizontal axis corresponds to the index of the LoRA experts.
@@ 5 Conclusions
In this paper, we first show that the task conflicts exists in vision language instruction tuning through extensive experiments. To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters. In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert. Comprehensive evaluations of MoCLE on 12 unseen tasks show the effectiveness of the proposed MoCLE method.
Acknowledgments.
We gratefully acknowledge support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
@@ Acknowledgments.
Appendix A Sources of Results in Table 1
For Flamingo-80B [2], Qwen-VL [3], MiniGPT-v2 [5], BLIP-2 [26], and InstructBLIP [11], the results are directly from their papers. For LLaVA and MiniGPT-4, we take the results from [5].
Appendix B Performance on LVLM Benchmarks
We evaluate MoCLE and the reproduced InstructBLIP on several LVLM benchmarks such as POPE [30], MME [13], and MM-Vet [59]. In details, given an image, POPE evaluates the object hallucination of LVLM by determining whether it could correctly respond to the question such as “Is there a broccoli in the image?”. MME and MM-Vet investigate various capacities of LVLM such as perception, cognition, etc. As shown in Table 8, incorporating MoCLE into InstructBLIP significantly reduces the object hallucination. This might result from the fact that questions w.r.t various object categories belong to different instruction clusters [^1] More sample instructions in the clusters are in Sec. H. and are routed to appropriate task experts in MoCLE. In addition, MoCLE also improves various abilities in LVLM as demonstrated in the performance gains in MME and MM-Vet.
Models POPE (adv.) POPE (rand.) POPE (pop.) MME (percep.) MM-Vet
InstructBLIP 77.6 84.3 78.2 1202.9 23.6
+ MoCLE 82.1 88.0 83.7 1222.6 26.0
Table 8: Performance of MoCLE on LVLM benchmarks. We report the F1-score, scores in perceptions, and the overall score for POPE, MME and MM-Vet, where a larger value corresponds to the better performance. POPE (adv.) indicates the adversarial objects splits of POPE. Similarly, POPE (rand.) and POPE (pop.) are for random and popular splits, respectively.
Appendix C Training Details
Following [11], we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP, (i.e., InstructBLIP (${}^{{\ddagger}}$) and the task experts in Sec. 1. We train those models with a maximum of 60K steps and a batch size of 128. The AdamW optimizer [23] is used, with $\beta_{1}$ as 0.9, $\beta_{2}$ as 0.999, and a weight decay as 0.05. We apply a linear warmup of the learning rate during the initial 1000 steps, increasing from ${10}^{-8}$ to ${10}^{-5}$, followed by a cosine decay with a minimum learning rate of 0.
Appendix D Weights of the Universal Experts
During training, if some training data obtain a very large weight on a task experts, such data tend to be very specific and might be less beneficial to other tasks. Hence, they get less weight on the universal expert. On the contrary, less specific (a.k.a, more general) data benefit more to other tasks and obtain larger weight on the universal expert. Therefore, the complementarity between the task experts and universal expert achieves good generalization in MoCLE.
@@ Appendix D Weights of the Universal Experts
Appendix E Datasets for Vision-language Instruction Tuning
Here we give a detailed description of the datasets used during instruction tuning. As shown in Table 9, we use 13 datasets in total. Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (Sec. F). For LLaVA-Instruct150K [33], we do not apply any task template as it is well formatted. For Web Capfilt, we use COCO+VG+CC+SBU (14M images) as introduced in [25]. In regard to which split is used, please refer to the original InstructBLIP [11] for more details. During training, we sample the datasets using the data balancing strategy proposed in [11].
@@ Appendix E Datasets for Vision-language Instruction Tuning
Appendix F Task Templates
Following [11], we use the same set of task templates for instruction tuning and zero-shot evaluation. Please refer to Tables 14 and 15 for training and evaluation templates.
Appendix G More Discussion
Limitation.
Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection [17, 27]) is appealing, which has shown more severe task conflicts with the conversation tasks [68]. Moreover, the applicability of MoCLE to image diffusion models [9, 14, 35, 29] is also a potential research direction, considering the ubiquity of negative transfer phenomenon.
@@ Appendix G More Discussion
Appendix H Case Studies
Conversations
In Table 10, we instruct the model to conduct a very difficult object counting task. The correct answer for this question is 63, which is quite hard for existing LVLMs. InstructBLIP failed to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
In Table G, the model is queried to recognize the character in the image. InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training. With MoCLE, the model can give correct results.
Similarly, in Table 13, we ask the model to describe the image in a detailed manner. InstructBLIP still mistakes this query as an image caption task and gives very short caption to this image. Differently, with MoCLE, the model correctly understands the “in details” in the instruction and provides sufficient details.
@@ Appendix H Case Studies
Sample Instructions in Clusters
In Table 16, we showcase some sample instructions assigned to different clusters. Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting. This justifies the use of a large number of instruction clusters. Moreover, in Table 5, we show that the model performs poorly with only a few clusters, i.e., $K=4$, and gets better as we increase $K$. With a few clusters, VQA, VQG, CONV and CAP tasks would occupy 4 clusters, leaving no more space for samples with different topics. As a result, the model might perform worse when mixing these topics.
@@
Template Group Task Template
CAP A short image caption:
A short image description:
A photo of
An image that shows
Write a short description for the image.
Write a description for the photo.
Provide a description of what is presented in the photo.
Briefly describe the content of the image.
Can you briefly explain what you see in the image?
Could you use a few words to describe what you perceive in the photo?
Please provide a short depiction of the picture.
Using language, provide a short account of the image.
Use a few words to illustrate what is happening in the picture.
VQA {Question}
Question: {Question}
Question A short answer to the question is
Q: {Question} A:
Question: {Question} Short answer:
Given the image, answer the following question with no more than three words. {Question}
Based on the image, respond to this question with a short answer: {Question}. Answer:
Use the provided image to answer the question: {Question} Provide your answer as short as possible:
What is the answer to the following question? “{Question}”
The question “{Question}” can be answered using the image. A short answer is
VQG Given the image, generate a question whose answer is: {Answer}. Question:
Based on the image, provide a question with the answer: {Answer}. Question:
Given the visual representation, create a question for which the answer is “{Answer}”.
From the image provided, craft a question that leads to the reply: {Answer}. Question:
Considering the picture, come up with a question where the answer is: {Answer}.
Taking the image into account, generate an question that has the answer: {Answer}. Question:
@@
Evaluation Dataset Instruction
GQA, VizWiz, iVQA, MSVD, MSRVTT Question: {} Short answer:
NoCaps, Flickr30k A short image description:
TextVQA OCR tokens: {}. Question: {} Short answer:
IconQA Question: {} Options: {}. Short answer:
ScienceQA Context: {} Question: {} Options: {}. Answer:
HatefulMemes This is an image with: “{}” written on it. Is it hateful? Answer:
VSR Based on the image, is this statement true or false? “{}” Answer:
@@
Cluster Instruction Samples Topics
1 “Q: what is being done to the food in the glass fronted box? A:” Food
@@
2 “Q: what sport is the cartoon dog playing? A:” Pet
@@
3 “what type dressing does this man favor?” Men
@@
4 “how many more animals need to be added to all of these to get the number ten?” Counting

### 22 Mar 2024 v4

@@ 4.5 Visualizations
Figure 3: (a) Clustering assignment of the training datasets when $K=64$. The labels on the vertical axis indicate the names of the datasets. Within each labeled block, multiple rows represent different subsets, each of which is associated with a unique task template. The horizontal axis denotes the cluster index to which these subsets are assigned. (b) t-SNE visualization of the instruction encoding. Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
4.5.1 Clustering.
@@
We first show the justification to represent the training data via their instructions. Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49]. After that, we visualize the data in Figure 2(b) via t-SNE [44]. According to Figure 2(b), we observe that (1) samples from different datasets but from the same task are grouped together. For example, coco_vqg, aok_vqg, and ok_vqg are all VQG (visual question generation) tasks (triangle markers) and reside on the left part of the figure. (2) samples from similar tasks are closed to each other. For example, coco_cap and textcaps both belong to the image captioning (CAP) task (small dots) and stay close to each other. Similarly, both the visual question answering (VQA, cross markers) task with datasets (e.g., aok_vqa, coco_vqa, ocrvqa, and ok_vqa) and conversation (CONV) task with datasets (e.g., llava_conversation, llava_detail, and llava_reason) involve answering user questions, which lie in the middle part of Figure 2(b), suggesting that instructions are good representatives of the training data.
@@
We then cluster all the instructions of the examples in the training data into 64 groups using $k$-means clustering. Figure 2(a) shows the cluster assignment of the training data. Here, each row in the heatmap denotes a subset of a dataset. The subset is obtained by applying the task template (Sec. 3.1) on the samples of the dataset. We observe (1) different subsets of the same datasets are assigned to similar clusters. For example, aok_vqa, coco_vqa, and ok_vqa are in the first several clusters, while aok_vqg, coco_vqg, and ok_vqg mainly occupy 4 other clusters. (2) datasets of similar tasks are assigned to common clusters. For example, llava_150k including llava_detail, llava_detail and llava_conversation and a series of VQA tasks share the first several clusters as they are to answer questions. Those evidences justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
4.5.2 Routing results.
@@
Figure 4 visualizes the routing decisions of our MoCLE and Sentence-MoLE. We obtain both results from one mixture of LoRA, i.e., one linear module in a layer. The routing results are aggregated by the subset of datasets similar to Figure 2(a). Shown in Figure 3(a), MoCLE could achieve task-level routing for the inputs. For example, datasets from VQA and VQG tasks are handled by expert 0 and 3 respectively. In addition, VQA expert (i.e., expert 0, which is mostly occupied by VQA task) is fundamental for most of the datasets. That is, most held-out datasets activate the VQA expert. This could be justified by the fact that most vision-language tasks take the form of question answering. However, this does not indicate that the entire model collapses to a single expert because knowledge from other tasks are learned by the universal expert and are leveraged during evaluation. Further, the remaining 3 experts are not dispensable as without them all tasks would clutter in a single expert, resulting in task conflicts. Instead, routing pattern of Sentence-MoLE in Figure 3(b) reveals little correlations between tasks/datasets and experts. That is, different datasets/tasks obtain similar routing decisions, still suffering from task conflicts.
@@
Figure 4: Routing decisions of one LoRA mixture for MoCLE and Sentence-MoLE. The setup of the vertical axis is similar to Figure 2(a) except that we also include the held-out tasks. They are separated by a dotted line on the vertical axis. The horizontal axis corresponds to the index of the LoRA experts.
@@ 5 Conclusions
In this paper, we first show that the task conflicts exists in vision language instruction tuning through extensive experiments. To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters. In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert. Comprehensive evaluations of MoCLE on 11 unseen tasks show the effectiveness of our MoCLE.
Acknowledgement
We gratefully acknowledge the support of MindSpore, CANN (Compute Architecture for Neural Networks) and Ascend AI Processor used for this research.
@@ Acknowledgement
Appendix 0.A Sources of Results in Table 1
For LLaVA (13B) and MiniGPT-4 (13B), the results on GQA, VSR, OKVQA and IconQA are from the Table 3 of [5], the results on POPE and A-OKVQA (Direct) are from Table 3 and 6 of [33], respectively. In particular, the performance of MiniGPT-4 (13B) on MME is from its official leaderboard  [^1] https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation. All the task performances of InstructBLIP (7B) are obtained from its reproduced version implemented in this paper.
Appendix 0.B Training Details
Following [14], we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP (7B) and the task experts in Sec. 1. We train those models with a maximum of 60K steps and a batch size of 128. The AdamW optimizer [26] is used, with $\beta_{1}$ as 0.9, $\beta_{2}$ as 0.999, and a weight decay as 0.05. We apply a linear warmup of the learning rate during the initial 1000 steps, increasing from ${10}^{-8}$ to ${10}^{-5}$, followed by a cosine decay with a minimum learning rate of 0.
Appendix 0.C Weights of the Universal Experts
During training, if some training data obtains a very large weight on a task expert, such data tend to be very specific and might be less beneficial to other tasks. Hence, they get less weight on the universal expert. On the contrary, less specific (a.k.a, more general) data benefit more to other tasks and obtain larger weight on the universal expert. Therefore, the complementarity between the task experts and universal expert achieves good generalization in MoCLE.
@@ Appendix 0.C Weights of the Universal Experts
Appendix 0.D Datasets for Vision-language Instruction Tuning
Here we give a detailed description of datasets used during instruction tuning. As shown in Table 9, we use 13 datasets in total. Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (see Appendix 0.E). For LLaVA-Instruct150K [38], we do not apply any task template as it has been well formatted. For Web Capfilt, we use COCO+VG+CC+SBU (14M images) as [29]. Regarding with which split is used, please refer to the InstructBLIP paper [14] for more details. During training, we sample datasets using the data balancing strategy as in [14].
Appendix 0.E Task Templates
Following [14], we use the same set of task templates for instruction tuning and held-in/out evaluation. Please refer to Tables 14 and 15 for training and evaluation templates.
Appendix 0.F Case Studies
0.F.1 Conversations
In Table 10, we instruct the model to conduct a very difficult object counting task. The correct answer for this question is 63, which is quite hard for existing LVLMs. InstructBLIP fails to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
In Table 11, the model is queried to recognize the character in the image. InstructBLIP performs not so well on this query possibly because OCR-related tasks conflict with other tasks during training. With MoCLE, the model can give correct results.
Similarly, in Table 13, we ask the model to describe the image in a detailed manner. InstructBLIP still mistakes this query as an image caption task and gives very short caption to this image. Instead, with MoCLE, the model correctly understands the “in details” in the instruction and provides sufficient details.
@@ 0.F.1 Conversations
0.F.2 Sample Instructions in Clusters
In Table 16, we showcase some sample instructions assigned to different clusters. Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting, justifying the usage of a large number of instruction clusters. Moreover, in Table 7, we show that the model performs poorly with only a few clusters, i.e., $K=4$, and gets better as we increase $K$. With a few clusters, VQA, VQG, CONV and CAP tasks will occupy 4 clusters, leaving no more space for samples with different topics. Thus, the model might perform worse when mixing these topics.
@@ 0.F.2 Sample Instructions in Clusters
Template Group
Task Template
CAP
“A short image caption:”
“A short image description:”
“A photo of”
“An image that shows”
“Write a short description for the image.”
“Write a description for the photo.”
“Provide a description of what is presented in the photo.”
“Briefly describe the content of the image.”
“Can you briefly explain what you see in the image?”
“Could you use a few words to describe what you perceive in the photo?”
“Please provide a short depiction of the picture.”
“Using language, provide a short account of the image.”
“Use a few words to illustrate what is happening in the picture.”
VQA
“{Question}”
“Question: {Question}”
“Question A short answer to the question is”
“Q: {Question} A:”
“Question: {Question} Short answer:”
“Given the image, answer the following question with no more than three words. {Question}”
“Based on the image, respond to this question with a short answer: {Question}. Answer:”
“Use the provided image to answer the question: {Question} Provide your answer as short as possible:”
“What is the answer to the following question? {Question}”
“The question {Question} can be answered using the image. A short answer is”
VQG
“Given the image, generate a question whose answer is: {Answer}. Question:”
“Based on the image, provide a question with the answer: {Answer}. Question:”
“Given the visual representation, create a question for which the answer is {Answer}.”
“From the image provided, craft a question that leads to the reply: {Answer}. Question:”
“Considering the picture, come up with a question where the answer is: {Answer}.”
“Taking the image into account, generate an question that has the answer: {Answer}. Question:”
@@
Evaluation Dataset
Instruction
GQA, iVQA,
“Question: {} Short answer:”
MSVD, MSRVTT,
OKVQA, VQAV2,
A-OKVQA(Direct)
A-OKVQA(MC),
“Question: {} Options: {}. Short answer:”
IconQA
NoCaps, Flickr30k
“A short image description:”
TextVQA
“OCR tokens: {}. Question: {} Short answer:”
ScienceQA
“Context: {} Question: {} Options: {}. Answer:”
HatefulMemes
“This is an image with: “{}” written on it. Is it hateful? Answer:”
VSR
“Based on the image, is this statement true or false? “{}” Answer:”
Visdual Dialog
@@
Cluster
Instruction Samples
Topics
1
“Q: what is being done to the food in the glass fronted box? A:”
Food
@@
2
“Q: what sport is the cartoon dog playing? A:”
Pet
@@
3
“what type dressing does this man favor?”
Men
@@
4
“how many more animals need to be added to all of these to get the number ten?”
Counting
@@
Appendix 0.G Limitation
Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA [18]) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection [20, 30]) is appealing, which has shown more severe task conflicts with the conversation tasks [74]. Moreover, the applicability of MoCLE to image diffusion models [9, 17, 39, 32] is also a potential research direction, considering the ubiquity of negative transfer phenomenon.